# Driving on Registers

### v1

@@ 1 Introduction
End-to-end (E2E) planning has emerged as a promising direction for autonomous driving (AD), offering a single pipeline that maps raw sensor data and ego state to driving decisions [hu2023uniad, paradrive]. Besides, by avoiding intermediate annotations such as 3D boxes, these methods reduce labeling cost. Among E2E approaches, trajectory-proposal methods, whether using a large pre-computed vocabulary [chen2024vadv2, li2024hydra, li2025hydramdppp, li2025generalized], or generating proposals on the fly [liao2025diffusiondrive, xing2025goalflow, guo2025ipad, feng2025rap], have shown particularly strong performance.
Methods predicting multiple possible trajectories and selecting between them naturally capture the uncertainty within navigation. As in model-based RL [moerland2023model], the ability to score becomes central: the scorer must reliably choose the best candidate using context encoded in the sensor features.
The sensor processing backbone producing the features that capture this context are thus key in E2E planning methods. These backbones typically dominate the parameter and FLOP count of E2E methods, often leveraging convolutional architectures like VoV-Net [lee2019energy] or large pre-trained networks such as Vision Transformers [Dosovitskiy2021vit] like EVA [fang2023eva], or DINO [oquab2024dinov2]. Such backbones output thousands of tokens per frame, which must be processed for hundreds of trajectories. This creates a major computational bottleneck that only worsens as resolution or sensor count increases.
The most common solution to reducing the bottleneck is to pool these features along the spatial dimensions. However, feature pooling enforces specific resolution requirements on sensor inputs, and treats all inputs as equally informative, performing the same averaging operation across all cameras. Inspired by works like [yu2024image], we ask just how many tokens are needed to represent a driving scene?
@@
Our method relies only on scoring annotations (rather than explicit 3D supervision) and achieves state-of-the-art results on NAVSIM-v1 [dauner2024navsim], NAVSIM-v2 [Cao2025navsim2], and the closed-loop HUGSIM benchmark [zhou2024hugsim]. Overall, our contributions are as follows:
@@ 2 Related Work
Token reduction is central to efficient ViTs, whose attention cost grows quadratically with sequence length. Simple solutions such as patch-group pooling are parameter-free but treat all tokens uniformly. Other training-free strategies include matching-based compression [bolya2023token]. Learned approaches range from Perceiver-IO’s latent queries [jaegle2022perceiverio] to ViT register tokens, originally introduced to fix attention sinks [darcet2023vision] and later used in compact generative models like TiTok [yu2024image]. Recent driving-focused works [ivanovic2025efficient, wang2025alpamayo] highlight the growing need for token reduction in real-time systems. To our knowledge, we are the first to repurpose ViT register tokens specifically for reducing visual tokens in E2E planning, enabling compact scene representations while retaining planning-critical context.
@@
End-to-end learning has become popular in autonomous driving since the proof of concept of pioneering works [hu2023uniad, jiang2023vad] like UniAD [hu2023uniad]. However, they still heavily rely on modular designs with different sub modules, such as detection, tracking and mapping, making it hard to deploy. With the introduction of more efficient (pseudo) closed-loop evaluation metrics [dauner2024navsim, Cao2025navsim2] or simulation [zhou2024hugsim], more recent methods [guo2025ipad, liao2025diffusiondrive, chitta2022transfuser, li2025recogdrive, chen2024vadv2] predict the planned trajectory or actions directly from sensor inputs, a step forward toward the fully end-to-end paradigm.
These E2E methods [guo2025ipad, liao2025diffusiondrive, li2025recogdrive] mostly rely on off-the-shelf CNN-based (e.g., ResNet-34, ResNet-50, V2-99) image encoders without putting further attention on the design of the perception stack. Downstream in the planning stack, transformers are more robustly explored. Transfuser [chitta2022transfuser] used transformers for LiDAR–image fusion; DriveTransformer [jia2025drivetransformer] unifies multiple tasks under one transformer; and iPad [guo2025ipad] exploits transformer residuals for iterative trajectory refinement. Innovation has concentrated on how to use mid-level sensor features rather than on stronger pretrained backbones. When ViTs [Dosovitskiy2021vit] are used, methods [feng2025rap, li2025generalized, yao2025drivesuprim, li2025ztrs] typically rely on large or huge variants without considering the computation costs. These models lean on attention to fuse heterogeneous inputs (ego state, images, LiDAR, commands), but often supplement it with costly 3D reasoning—BEV projections, deformable cross-attention, or LiDAR supervision, raising annotation and compute demands. In this context, we design a simple query-based transformer architecture, built on the much smaller ViT variants, avoiding the high cost of complex intermediate representation, while keeping superior E2E driving performance.
@@
Trajectory-proposal based planning has become a compelling direction for E2E driving, introduced in Hydra-MDP [li2025hydramdppp]. Producing many possible futures and selecting one forces the model to address multimodality head-on. This shifts importance toward the scoring module, which must reliably choose the best candidate [liao2025diffusiondrive]. A recent state-of-the-art work on NAVSIM-v2, GTRS [li2025generalized] showed that learning a sufficiently strong scorer using a ViT backbone, paired with a large trajectory vocabulary, can solve complex scenarios. We introduce the importance of separate scoring and trajectory generation pathways in our E2E model.
@@ 3.1 A Simple Design
We design a simple and efficient transformer based architecture for planning in autonomous driving. The overall pipeline is presented in Figure 1. It is composed of three modules: an encoder for perception and two decoders for trajectory estimation and scoring. The model follows a classical transformer encoder-decoder architecture [vaswani2017attention] without any complex intermediate representation.
The encoder for perception (see subsection 3.2) is a vision transformer [Dosovitskiy2021vit] applied per camera. In order to compress the feature map to small set of tokens, we introduce additional per-camera registers, finetuned with the backbone to encode the perceptual information. We group together all per-camera register tokens at the output of the ViT to form the scene tokens.
@@ 3.2 Perception Encoder
This compression into a small set of camera tokens is close in spirit to Perceiver approaches [jaegle2021perceiver, jaegle2022perceiverio], with the noticeable difference that these approaches use cross-attentions for compression. Setting up such a mechanism in the encoder would require changes in the ViT architecture.
@@ 3.3 Trajectories
All decoders use the architecture depicted in Figure 2b, which consists of a vanilla transformer decoder [vaswani2017attention]: a stack of $k$ transformer blocks, each made of a self-attention layer, followed by a cross-attention to the scene tokens, and a feed-forward network (FFN), all with residual connections.
@@
The trajectories $\tau_{i}$} are learned using a Winner-Takes-All (WTA) approach [guzman2012multiple] or, equivalently, the minimum-over-n (MoN) loss [guo2025ipad]. Given a reference human trajectory $\hat{\tau}$ of duration $T$ and consisting of $n_{p}$ poses, only the closest predicted trajectory is supervised, which allows the model to produce diverse candidate trajectories:
@@ 3.4 Scores
Finally we predict the six score components [dauner2024navsim] using a dedicated MLP for each score.
@@
At inference, we interpret our scoring network as a reward function and our full pipeline as a driving policy which is conditioned on a specific behavior profile encapsulated in the set of weights applied to the scoring outputs. Borrowing from Offline-RL literature such as CtRL-Sim [rowe2024ctrl], we can thus condition our final scoring output on a driving behavior by modifying the values of our $\lambda_{c}$, and selecting trajectories which maximize the score computed with this new set of weights. For example, this can encourage trajectories which make maximal estimated progress.
@@ 4 Experiments
As input, we use 4 cameras (front, front left, front right and back). The perception module is a DINOv2 ViT-S [oquab2024dinov2] encoder, LoRA finetuned [hu2022lora] (rank 32) following [barin2024robust]. By default, we add 16 registers per camera. The decoders are 4-layer transformers with an inner dimension of 256. The feed- forward network has a dilation factor of 4. The camera registers as well as the initial trajectory tokens are randomly initialized with normal distribution $\mathcal{N}(0,10^{-6})$. The model is trained on the navtrain split for 10 epochs, with learning rate $2\times 10^{-4}$ and cosine annealing scheduling. For the ablation studies, if not described otherwise, we train our model with this common base architecture. All ablation model scores are computed on the navval split. All models were trained on 4 NVIDIA A100 GPUs. The DrivoR model is roughly 40M parameters, significantly below comparable works.
@@ 4.1 Benchmarks
NAVSIM-v1 [dauner2024navsim] is a dataset built out of nuPlan [caesar2021nuplan] as a subset of OpenScene [contributors2023openscene]. As opposed to previous benchmark such as nuScenes [nuscenes], mostly formulating driving quality as a measure of similarity to the expert human trajectory, NAVSIM-v1 introduces metrics inspired from closed-loop simulation. The main metric, Predictive Driver Model Score (PDMS), is an aggregation of penalties, e.g., collisions with other, non-reactive, agents, staying on-road and in the correct direction, and avoiding near-misses with other agents. These scores are combined with quality-related scores such as the comfort and progress. Progress is measured as a comparison to the centerline-progress of a PDM agent [dauner2023pdm] given access to complete ground-truth information. For training, we use default PDMS weights defined in the benchmark [dauner2024navsim].
The results obtained on the NAVSIM-v1 benchmark are presented in Table 1. DrivoR outperforms all other methods on NAVSIM-v1, and nears human-level performance. We highlight the comparison to RAP [feng2025rap], which we consider an orthogonal work: the large quantity of rasterized data introduced in the paper may be used in any method.
@@
PDM‑Closed [dauner2023pdm] PMLR’23 94.6 99.8 89.9 86.9 99.9 89.1
Human driver [dauner2024navsim] NeurIPS’24 100 100 100 99.9 87.5 94.8
RAP-DINO [feng2025rap] arXiv’25 99.1 98.9 96.7 100 90.3 93.8
Ego‑stat. MLP [dauner2024navsim] NeurIPS’24 93.0 77.3 83.6 100 62.8 65.6
UniVLA [wang2025unified] arXiv’25 96.9 91.1 91.7 96.7 76.8 81.7
DrivingGPT [chen2024drivinggpt] ICCV’24 98.9 90.7 94.9 95.6 79.7 82.4
UniAD [hu2023uniad] CVPR’23 97.8 91.9 92.9 100 78.8 83.4
LTF [chitta2022transfuser] TPAMI’22 97.4 92.8 92.4 100 79.0 83.8
PARA‑Drive [paradrive] CVPR’24 97.9 92.4 93.0 99.8 79.3 84.0
DriveX-S [shi2025drivex] ICCV’25 97.5 94.0 93.0 100 79.7 84.5
World4Drive [zheng2025world4drive] ICCV’25 97.4 94.3 92.8 100 79.9 85.1
DRAMA [yuan2024drama] ISRR’24 98.0 93.1 94.8 100 80.1 85.5
VAD-v2 [chen2024vadv2] arXiv’24 98.1 94.8 94.3 100 80.6 86.2
PRIX [wozniak2025prix] arXiv’25 98.1 96.3 94.1 100 82.3 87.8
DiffusionDrive [liao2025diffusiondrive] CVPR’25 98.2 96.2 94.7 100 82.2 88.1
DIVER [song2025breaking] arXiv’25 98.5 96.5 94.9 100 82.6 88.3
AutoVLA [zhou2025autovla] NeurIPS’25 98.4 95.6 98.0 99.9 81.9 89.1
DriveVLA-W0 [li2025drivevla] arXiv’25 98.7 99.1 95.3 99.3 83.3 90.2
ReCogDrive [li2025recogdrive] arXiv’25 97.9 97.3 94.9 100 87.3 90.8
Hydra-MDP++ [li2025hydramdppp] arXiv’25 98.6 98.6 95.1 100 85.7 91.0
iPad [guo2025ipad] arXiv’25 98.6 98.3 94.9 100 88.0 91.7
Centaur [sima2025centaur] arXiv’25 99.5 98.9 98.0 100 85.9 92.6
DriveSuprim [yao2025drivesuprim] arXiv’25 98.6 98.6 95.5 100 91.3 93.5
DrivoR 99.0 98.9 96.7 100 90.0 93.7
@@
VAD [jiang2023vad] 51.3 31.1 25.3 26.5 31.4 36.3 09.5 08.0 11.5 13.4
LTF [chitta2022transfuser] 67.8 35.1 26.2 40.5 38.9 58.9 18.0 09.8 25.9 23.7
UniAD [hu2023uniad] 78.4 60.5 33.6 17.8 45.9 64.9 45.8 20.6 06.6 32.7
@@
Table 2: Photorealistic closed-loop evaluation on HUGSIM [zhou2024hugsim]. Zero-shot generalization using the DrivoR model from the NAVSIM-v1 evaluation. Scores are per difficulty and overall average road completion (RC) and HD-Score, higher always better.
@@
RAP-DINO (ViT-H) [feng2025rap] 97.1 94.4 98.8 99.8 83.9 96.9 94.7 96.4 66.2 83.2 83.9 87.4 98.0 86.9 80.4 52.3 95.2 52.4 39.6
GTRS-D (V2-99) [li2025generalized] 98.9 96.2 99.4 99.3 72.9 98.9 95.1 96.9 39.1 91.2 89.4 94.4 98.8 69.5 90.0 54.3 94.0 48.7 45.0
GTRS-A (V2-99) [li2025generalized] 98.9 95.1 99.1 99.6 76.2 99.1 94.9 97.6 54.2 88.1 88.8 89.3 98.9 98.9 85.9 53.7 96.8 56.9 45.4
@@
NAVSIM-v2 [Cao2025navsim2] builds on NAVSIM-v1 with the objective of closing the gap with closed-loop, simulator driven, benchmarks. NAVSIM-v2 introduces a second stage of evaluation, where novel variations of scenes are generated via Gaussian Splatting. These novel scenes consist of perturbations of the ego vehicle status, i.e., shifts and rotations, forcing the model to generalize outside of the training distribution. NAVSIM-v2 is scored using an extended version of the score from NAVSIM-v1, termed the EPDMS. We present the results on the navhard-two-stage split of NAVSIM-v2 in table Table 3, where DrivoR outperforms all existing works. We note that the results in Table 3 were computed after an official bug fix, and thus do not include the largest GTRS-Dense model. Results on NAVSIM-2 before the fix are included in the supplementary material. We use warmup-two-stage111warmup-two-stage intersects with navhard-two-stage, after request, benchmark authors validated its use as validation set. as validation set for NAVSIM-v2 containing 7 scenes.
@@
For the closed-loop evaluation, we use HUGSIM [zhou2024hugsim], a benchmark with scenarios adapted from the original scenes of KITTI-360 [kitti360], nuScenes [nuscenes], PandaSet [pandaset], and Waymo [waymo]. These scenarios are reconstructed as photorealistic 3D environments in which an E2E model controls the ego agent to navigate within each scenario. The planner perceives the 3D environment through RGB cameras whose viewpoints are determined by the ego agent’s position in the scenario.

### v2

@@ 1 Introduction
End-to-end (E2E) planning has emerged as a promising direction for autonomous driving (AD), offering a single pipeline that maps raw sensor data and ego state to driving decisions [19, 49]. Besides, by avoiding intermediate annotations such as 3D boxes, these methods reduce labeling cost. Among E2E approaches, trajectory-proposal methods, whether using a large pre-computed vocabulary [6, 33, 29, 35], or generating proposals on the fly [36, 52, 16, 15], have shown particularly strong performance.
Methods predicting multiple possible trajectories and selecting between them naturally capture the uncertainty within navigation. As in model-based RL [38], the ability to score becomes central: the scorer must reliably choose the best candidate using context encoded in the sensor features.
The sensor processing backbone producing the features that capture this context are thus key in E2E planning methods. These backbones typically dominate the parameter and FLOP count of E2E methods, often leveraging convolutional architectures like VoV-Net [27] or large pre-trained networks such as Vision Transformers [13] like EVA [14], or DINO [39]. Such backbones output thousands of tokens per frame, which must be processed for hundreds of trajectories. This creates a major computational bottleneck that only worsens as resolution or sensor count increases.
The most common solution to reducing the bottleneck is to pool these features along the spatial dimensions. However, feature pooling enforces specific resolution requirements on sensor inputs, and treats all inputs as equally informative, performing the same averaging operation across all cameras. Inspired by works like [54], we ask just how many tokens are needed to represent a driving scene?
@@
Our method relies only on scoring annotations (rather than explicit 3D supervision) and achieves state-of-the-art results on NAVSIM-v1 [12], NAVSIM-v2 [5], and the closed-loop HUGSIM benchmark [60]. Overall, our contributions are as follows:
@@ 2 Related Work
Token reduction is central to efficient ViTs, whose attention cost grows quadratically with sequence length. Simple solutions such as patch-group pooling are parameter-free but treat all tokens uniformly. Other training-free strategies include matching-based compression [2]. Learned approaches range from Perceiver-IO’s latent queries [21] to ViT register tokens, originally introduced to fix attention sinks [10] and later used in compact generative models like TiTok [54]. Recent driving-focused works [20, 47] highlight the growing need for token reduction in real-time systems. To our knowledge, we are the first to repurpose ViT register tokens specifically for reducing visual tokens in E2E planning, enabling compact scene representations while retaining planning-critical context.
@@
End-to-end learning has become popular in autonomous driving since the proof of concept of pioneering works [19, 24] like UniAD [19]. However, they still heavily rely on modular designs with different sub modules, such as detection, tracking and mapping, making it hard to deploy. With the introduction of more efficient (pseudo) closed-loop evaluation metrics [12, 5] or simulation [60], more recent methods [16, 36, 8, 32, 6] predict the planned trajectory or actions directly from sensor inputs, a step forward toward the fully end-to-end paradigm.
These E2E methods [16, 36, 32] mostly rely on off-the-shelf CNN-based (e.g., ResNet-34, ResNet-50, V2-99) image encoders without putting further attention on the design of the perception stack. Downstream in the planning stack, transformers are more robustly explored. Transfuser [8] uses transformers for LiDAR–image fusion; DriveTransformer [23] unifies multiple tasks under one transformer; and iPad [16] exploits transformer residuals for iterative trajectory refinement. Innovation has concentrated on how to use mid-level sensor features rather than on stronger pretrained backbones. When ViTs [13] are used, methods [15, 35, 53, 34] typically rely on large or huge variants without considering the computation costs. These models lean on attention to fuse heterogeneous inputs (ego state, images, LiDAR, commands), but often supplement it with costly 3D reasoning-BEV projections, deformable cross-attention, or LiDAR supervision, raising annotation and compute demands. In this context, we design a simple query-based transformer architecture, built on the much smaller ViT variants, avoiding the high cost of complex intermediate representation, while keeping superior E2E driving performance.
@@
Trajectory-proposal based planning has become a compelling direction for E2E driving, introduced in Hydra-MDP [29]. Producing many possible futures and selecting one forces the model to address multimodality head-on. This shifts importance toward the scoring module, which must reliably choose the best candidate [36]. A recent state-of-the-art work on NAVSIM-v2, GTRS [35] showed that learning a sufficiently strong scorer using a ViT backbone, paired with a large trajectory vocabulary, can solve complex scenarios. We introduce the importance of separate scoring and trajectory generation pathways in our E2E model.
@@ 3.1 A Simple Design
We design a simple and efficient transformer based architecture for planning in autonomous driving. The overall pipeline is presented in Figure 1. It is composed of three modules: an encoder for perception and two decoders for trajectory estimation and scoring. The model follows a classical transformer encoder-decoder architecture [46] without any complex intermediate representation.
The encoder for perception (see subsection 3.2) is a vision transformer [13] applied per camera. In order to compress the feature map to small set of tokens, we introduce additional per-camera registers, finetuned with the backbone to encode the perceptual information. We group together all per-camera register tokens at the output of the ViT to form the scene tokens.
@@ 3.2 Perception Encoder
This compression into a small set of camera tokens is close in spirit to Perceiver approaches [22, 21], with the noticeable difference that these approaches use cross-attentions for compression. Setting up such a mechanism in the encoder would require changes in the ViT architecture.
@@ 3.3 Trajectories
All decoders use the architecture depicted in Figure 2b, which consists of a vanilla transformer decoder [46]: a stack of $k$ transformer blocks, each made of a self-attention layer, followed by a cross-attention to the scene tokens, and a feed-forward network (FFN), all with residual connections.
@@
The trajectories $\tau_{i}$} are learned using a Winner-Takes-All (WTA) approach [17] or, equivalently, the minimum-over-n (MoN) loss [16]. Given a reference human trajectory $\hat{\tau}$ of duration $T$ and consisting of $n_{p}$ poses, only the closest predicted trajectory is supervised, which allows the model to produce diverse candidate trajectories:
@@ 3.4 Scores
Finally we predict the six score components [12] using a dedicated MLP for each score.
@@
At inference, we interpret our scoring network as a reward function and our full pipeline as a driving policy which is conditioned on a specific behavior profile encapsulated in the set of weights applied to the scoring outputs. Borrowing from Offline-RL literature such as CtRL-Sim [40], we can thus condition our final scoring output on a driving behavior by modifying the values of our $\lambda_{c}$, and selecting trajectories which maximize the score computed with this new set of weights. For example, this can encourage trajectories which make maximal estimated progress.
@@ 4 Experiments
As input, we use 4 cameras (front, front left, front right and back). The perception module is a DINOv2 ViT-S [39] encoder, LoRA finetuned [18] (rank 32) following [1]. By default, we add 16 registers per camera. The decoders are 4-layer transformers with an inner dimension of 256. The feed- forward network has a dilation factor of 4. The camera registers as well as the initial trajectory tokens are randomly initialized with normal distribution $\mathcal{N}(0,10^{-6})$. The model is trained on the navtrain split for 10 epochs, with learning rate $2\times 10^{-4}$ and cosine annealing scheduling. For the ablation studies, if not described otherwise, we train our model with this common base architecture. All ablation model scores are computed on the navval split. All models were trained on 4 NVIDIA A100 GPUs. The DrivoR model is roughly 40M parameters, significantly below comparable works.
@@ 4.1 Benchmarks
NAVSIM-v1 [12] is a dataset built out of nuPlan [4] as a subset of OpenScene [9]. As opposed to previous benchmark such as nuScenes [3], mostly formulating driving quality as a measure of similarity to the expert human trajectory, NAVSIM-v1 introduces metrics inspired from closed-loop simulation. The main metric, Predictive Driver Model Score (PDMS), is an aggregation of penalties, e.g., collisions with other, non-reactive, agents, staying on-road and in the correct direction, and avoiding near-misses with other agents. These scores are combined with quality-related scores such as the comfort and progress. Progress is measured as a comparison to the centerline-progress of a PDM agent [11] given access to complete ground-truth information. For training, we use default PDMS weights defined in the benchmark [12].
The results obtained on the NAVSIM-v1 benchmark are presented in Table 1. DrivoR outperforms all other methods on NAVSIM-v1, and nears human-level performance. We highlight the comparison to RAP [15], which we consider an orthogonal work: the large quantity of rasterized data introduced in the paper may be used in any method.
@@
PDM‑Closed [11] PMLR’23 94.6 99.8 89.9 86.9 99.9 89.1
Human driver [12] NeurIPS’24 100 100 100 99.9 87.5 94.8
RAP-DINO [15] arXiv’25 99.1 98.9 96.7 100 90.3 93.8
Ego‑stat. MLP [12] NeurIPS’24 93.0 77.3 83.6 100 62.8 65.6
UniVLA [48] arXiv’25 96.9 91.1 91.7 96.7 76.8 81.7
DrivingGPT [7] ICCV’24 98.9 90.7 94.9 95.6 79.7 82.4
UniAD [19] CVPR’23 97.8 91.9 92.9 100 78.8 83.4
LTF [8] TPAMI’22 97.4 92.8 92.4 100 79.0 83.8
PARA‑Drive [49] CVPR’24 97.9 92.4 93.0 99.8 79.3 84.0
DriveX-S [41] ICCV’25 97.5 94.0 93.0 100 79.7 84.5
World4Drive [58] ICCV’25 97.4 94.3 92.8 100 79.9 85.1
DRAMA [56] ISRR’24 98.0 93.1 94.8 100 80.1 85.5
VAD-v2 [6] arXiv’24 98.1 94.8 94.3 100 80.6 86.2
PRIX [50] arXiv’25 98.1 96.3 94.1 100 82.3 87.8
DiffusionDrive [36] CVPR’25 98.2 96.2 94.7 100 82.2 88.1
DIVER [43] arXiv’25 98.5 96.5 94.9 100 82.6 88.3
AutoVLA [61] NeurIPS’25 98.4 95.6 98.0 99.9 81.9 89.1
DriveVLA-W0 [30] arXiv’25 98.7 99.1 95.3 99.3 83.3 90.2
ReCogDrive [32] arXiv’25 97.9 97.3 94.9 100 87.3 90.8
Hydra-MDP++ [29] arXiv’25 98.6 98.6 95.1 100 85.7 91.0
iPad [16] arXiv’25 98.6 98.3 94.9 100 88.0 91.7
Centaur [42] arXiv’25 99.5 98.9 98.0 100 85.9 92.6
DriveSuprim [53] arXiv’25 98.6 98.6 95.5 100 91.3 93.5
DrivoR (train) 98.9 98.3 96.2 100 89.1 93.1
DrivoR (trainval) 99.0 98.9 96.7 100 90.0 93.7
DrivoR (+65k SimScale data) 99.1 99.0 96.9 100 90.3 94.0
DrivoR (+134k SimScale data) 99.1 99.2 96.9 100 91.6 94.6
@@
VAD [24] 51.3 31.1 25.3 26.5 31.4 36.3 09.5 08.0 11.5 13.4
LTF [8] 67.8 35.1 26.2 40.5 38.9 58.9 18.0 09.8 25.9 23.7
UniAD [19] 78.4 60.5 33.6 17.8 45.9 64.9 45.8 20.6 06.6 32.7
@@
Table 2: Photorealistic closed-loop evaluation on HUGSIM [60]. Zero-shot generalization using the DrivoR model from the NAVSIM-v1 evaluation. Scores are per difficulty and overall average road completion (RC) and HD-Score, higher always better.
@@
RAP-DINO (ViT-H) [15] 97.1 94.4 98.8 99.8 83.9 96.9 94.7 96.4 66.2 83.2 83.9 87.4 98.0 86.9 80.4 52.3 95.2 52.4 39.6
GTRS-D (V2-99) [35] 98.9 96.2 99.4 99.3 72.9 98.9 95.1 96.9 39.1 91.2 89.4 94.4 98.8 69.5 90.0 54.3 94.0 48.7 45.0
GTRS-A (V2-99) [35] 98.9 95.1 99.1 99.6 76.2 99.1 94.9 97.6 54.2 88.1 88.8 89.3 98.9 98.9 85.9 53.7 96.8 56.9 45.4
@@
DrivoR (+65k SimScale data, ViT-S) 98.9 97.3 99.2 99.6 77.7 99.1 95.3 97.6 68.4 92.3 92.2 97.0 99.0 72.1 90.3 56.3 97.1 38.8 52.3
SimScale (+185k SimScale data, V2-99) [45] 99.6 99.16 99.9 100 69.6 99.6 95.8 95.6 28.4 94.5 94.2 95.8 99.2 75.8 92.8 60.1 96.1 43.2 53.2
DrivoR (+134k SimScale data, ViT-S) 99.1 98.2 99.3 99.8 75.4 98.7 94.9 97.6 70.2 92.3 91.6 97.3 99.1 75.7 90.6 56.1 98.4 44.7 54.6
@@
NAVSIM-v2 [5] builds on NAVSIM-v1 with the objective of closing the gap with closed-loop, simulator driven, benchmarks. NAVSIM-v2 introduces a second stage of evaluation, where novel variations of scenes are generated via Gaussian Splatting. These novel scenes consist of perturbations of the ego vehicle status, i.e., shifts and rotations, forcing the model to generalize outside of the training distribution. NAVSIM-v2 is scored using an extended version of the score from NAVSIM-v1, termed the EPDMS. We present the results on the navhard-two-stage split of NAVSIM-v2 in table Table 3, where DrivoR outperforms all existing works. We note that the results in Table 3 were computed after an official bug fix, and thus do not include the largest GTRS-Dense model. Results on NAVSIM-2 before the fix are included in the supplementary material. We use warmup-two-stage111warmup-two-stage intersects with navhard-two-stage, after request, benchmark authors validated its use as validation set. as validation set for NAVSIM-v2 containing 7 scenes.
@@
For the closed-loop evaluation, we use HUGSIM [60], a benchmark with scenarios adapted from the original scenes of KITTI-360 [37], nuScenes [3], PandaSet [51], and Waymo [44]. These scenarios are reconstructed as photorealistic 3D environments in which an E2E model controls the ego agent to navigate within each scenario. The planner perceives the 3D environment through RGB cameras whose viewpoints are determined by the ego agent’s position in the scenario.
@@
(a) NAVSIM-v1 (PDMS)
(b) NAVSIM-v2 (EPDMS)
Figure 3: DrivoR Scalability. Scaling DrivoR (blue) against current SOTA baselines (orange). (a) On the NAVSIM-v1 task, DrivoR surpasses both the RAP [15] VIT-H baseline. (b) On the NAVSIM-v2 task, DrivoR achieves a new state-of-the-art score of 54.6 EPDMS.
Scaling DrivoR.
With its simple design using vanilla transformers and registers, we study the scalability of DrivoR. We leverage synthetic data from [45] with pseudo annotations from PDM-Closed [11]. Concretely, we extract 65k and 134k annotations from [45] and jointly train with navtrain, respectively. We show in Figure 3 that DrivoR consistently improves the NAVSIM performance with the increased data. Moreover, DrivoR sets a new state-of-the-art performance with less data, compared to RAP [15] using 500k+ synthetic data and SimScale [45] with 185k extra data. More detailed performance scores are provided in Table 1 and Table 3.